# Finding all solutions of qKZ equations in characteristic $p$

### v1

@@ 2.5. KZ equations as a limit of qKZ equations
(2.17) $f(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})\,=\,K_{a}(z\>;\kappa)\,f(z),\qquad a=1,\dots,n.\ignorespaces$
@@
(2.18) $g(w_{1},\dots,w_{a}-\alpha\kappa,\dots,w_{n};\alpha)\,=\,K_{a}(w/\alpha\>;\kappa)\,g(w;\alpha),\qquad a=1,\dots,n,\ignorespaces$
@@
$K_{a}(w/\alpha;\kappa)$ $=$ $R^{(a,a-1)}(w_{a}-w_{a-1}-\alpha\kappa;\alpha)\,\dots\,R^{(a,1)}(w_{a}-w_{1}-\alpha\kappa;\alpha)$
$\times$ $R^{(a,n)}(w_{a}-w_{n};\alpha)\,\dots\,R^{(a,a+1)}(w_{a}\mskip-0.99998mu-w_{a+1};\alpha)\,,$
@@
$g-\alpha\kappa\frac{\partial g}{\partial w_{a}}+\mathcal{O}(\alpha^{2})=(1-\alpha H_{a}(w)+\mathcal{O}(\alpha^{2}))g.$
In the limit $\alpha\to 0$, we obtain the the KZ differential equations
(2.20) $\kappa\,\frac{\partial g}{\partial w_{a}}(w)=H_{a}(w)g(w)\,.$
@@
(2.21) $C_{a}^{\operatorname{KZ}}(z;\kappa)=\lim_{\alpha\to 0}\hat{C}_{a}(z_{1}/\alpha,\dots,z_{n}/\alpha;\kappa)/\alpha^{p}.$
@@
(2.22) $C_{a}^{\operatorname{KZ}}(w;\kappa)=\lim_{\alpha\to 0}(C_{a,\alpha}(w;\kappa)-1)/\alpha^{p}.$
@@
$(\nabla_{a,\alpha}g)(w)=g-\alpha H_{a}(w)g+\alpha\kappa\frac{\partial g}{\partial w_{a}}+\mathcal{O}(\alpha^{2})$
@@
$(\nabla_{a,\alpha}-1)g=\alpha\nabla^{\operatorname{KZ}}_{a}g+\mathcal{O}(\alpha^{2}).$
@@
$C_{a,\alpha}(w;\kappa)-1=(\nabla_{a,\alpha})^{p}-1=(\nabla_{a,\alpha}-1)^{p}=\alpha^{p}\left(\nabla^{\operatorname{KZ}}_{a}\right)^{p}+\mathcal{O}(\alpha^{p+1}).$
@@
The degree of the reduced $p$-curvature operator $\tilde{C}_{a}(z;\kappa)$ as a polynomial in $z$ is not greater than $(n-2)p$. Moreover, we have
(2.23) $\tilde{C}_{a}(z;\kappa)=\tilde{C}_{a}^{\operatorname{KZ}}(z;\kappa)+f_{1}(z)$
where $f_{1}(z)$ is a polynomial in $z$ of degree less than $(n-2)p$.
@@
The first statement of the corollary follows from the fact that the denominator of the reduced $p$-curvature operator $\hat{C}_{a}(z;\kappa)$ is of degree $(n-1)p$. Hence the numerator is of degree not greater than $(n-2)p$ for the limit in (2.22) to exist. The second statement is clear. ∎
2.6. Limit of solutions
Let $f(z_{1},\dots,z_{n})$ be a polynomial solution of the qKZ equations (2.4). Let $\deg f(z)=d$ and
$f(z)=f_{0}(z)+f_{1}(z),$
where $f_{0}(z)$ is a homogeneous polynomial of degree $d$ and $f_{1}(z)$ is a polynomial of degree less than $d$.
@@ 2.6. Limit of solutions
Let $\alpha$ be a formal parameter. Define $g(w_{1},\dots,w_{n};\alpha)=\alpha^{d}f(w_{1}/\alpha,\dots,w_{n}/\alpha)$. Then
$g(w_{1},\dots,w_{n};\alpha)=f_{0}(w)+\alpha g_{1}(w;\alpha)$
where $g_{1}(w;\alpha)=\alpha^{d-1}f_{1}(w/\alpha)$ is a polynomial in $w$ and $\alpha$. We have
(2.24) $g(w_{1},\dots,w_{a}-\alpha\kappa,\dots,w_{n};\alpha)\,=\,K_{a}(w/\alpha\>;\kappa)\,g(w;\alpha),\qquad a=1,\dots,n.\ignorespaces$
@@
$n_{a}(w;\alpha)$ $=$ $(w_{a}-w_{a-1}-\alpha\kappa-\alpha P^{(a,a-1)})\dots(w_{a}-w_{a+1}-\alpha P^{(a,a+1)})$
@@
(2.25) $d_{a}(w;\alpha)\,g(w_{1},\dots,w_{a}-\alpha\kappa,\dots,w_{n};\alpha)\,=\,n_{a}(w;\alpha)\,g(w;\alpha)$
in the variables $w,\alpha$. Equation (2.25) gives an equation in $w$ for every fixed power of $\alpha$ in (2.25). The equation corresponding to the first power of $\alpha$ after division by $d_{a}(w;0)$ becomes
(2.26) $\kappa\,\frac{\partial f_{0}}{\partial w_{a}}(w)=\sum_{j\neq a}\frac{P^{(a,j)}-1}{w_{a}-w_{j}}\,f_{0}(w)\,.$
2.7. Dual qKZ equations
Denote $W=L^{\otimes n}$. Let $W^{*}$ be the dual space of $W$, and $\langle\,,\,\rangle:W^{*}\otimes W\to{\mathbb{K}}$ the canonical pairing. Let $K_{a}^{*}(z;\kappa):W^{*}\to W^{*}$ be the operators dual to the operators $K_{a}(z;\kappa)$. Denote $\tilde{K}_{a}(z;\kappa)=(K_{a}^{*}(z;\kappa))^{-1}$.
@@ 2.7. Dual qKZ equations
$\tilde{K}_{a}(z_{1},\dots,z_{b}-\kappa,\dots,z_{n};\kappa)\,\tilde{K}_{b}(z;\kappa)=\tilde{K}_{b}(z_{1},\dots,z_{a}-\kappa\dots,z_{n};\kappa)\,\tilde{K}_{a}(z;\kappa)$
@@
(2.27) $\langle x,y\rangle=\left\langle\tilde{K}_{a}(z;\kappa)x,K_{a}(z;\kappa)y\right\rangle$
for all all $a$ and $x\in W^{*},\,y\in W$.
@@
(2.28) $\tilde{s}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})\,=\,\tilde{K}_{a}(z\>;\kappa)\,\tilde{s}(z),\qquad a=1,\dots,n,\ignorespaces$
@@
(2.29) $\langle\tilde{s}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n}),s(z_{1},\dots,z_{a}-\kappa,\dots,z_{n},)\rangle=\langle\tilde{s}(z),s(z)\rangle.$
The set of vectors $\{v_{I}\mid I\subset\{1,\dots,n\}\}$ is a basis of $W$. Define the nondegenerate symmetric bilinear form $S$ on $W$ by the formula
(2.30) $S(v_{I},v_{J})=\delta_{I,J}\,.$
The form is called the (tensor) Shapovalov form. We identify $W^{*}$ and $W$ with the help of the Shapovalov form.
@@
(2.31) $\tilde{K}_{a}(z;\kappa)=R^{(a,a-1)}(-z_{a}+z_{a-1}+\kappa)\,\dots\,R^{(a,1)}(-z_{a}+z_{1}+\kappa)$
×$\hskip 120.0pt\times\ R^{(a,n)}(-z_{a}+z_{n})\,\dots\,R^{(a,a+1)}(-z_{a}\mskip-0.99998mu+z_{a+1})\,.$
@@
(2.32) $\tilde{K}_{a}(z_{1},\dots,z_{n};\kappa)$ $=$ $K_{a}(-z_{1},\dots,-z_{n};-\kappa),$
@@
(2.33) $\tilde{s}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})\,=\,K_{a}(-z_{1},\dots,-z_{n};-\kappa)\,s(z),\qquad a=1,\dots,n.\ignorespaces$
@@
Let $s(z)$ be a solution of the qKZ equations (2.4) with step $\kappa$ and $\tilde{s}(z)$ a solution of the qKZ equations (2.4) with step $-\kappa$. Then the function $S(\tilde{s}(-z),s(z))$ is $\kappa$-periodic with respect to every variable $z_{a}$ ,
(2.34) $\quad S(\tilde{s}(-z_{1},\dots,-(z_{a}-\kappa),\dots,-z_{n}),s(z_{1},\dots,z_{a}-\kappa,\dots,z_{n}))=S(\tilde{s}(-z),s(z)).$
@@
3. Pochhammer polynomials
For $\kappa\in{\mathbb{K}}$ and $m$ a positive integer, define Pochhammer polynomial $(t;\kappa)_{m}\in{\mathbb{K}}[t]$ by the formula
(3.1) $(t;\kappa)_{m}=\prod_{i=1}^{m}(t-(i-1)\kappa),$
see [FRF]. We have
@@ 3. Pochhammer polynomials
(3.2) $(t+z;\kappa)_{m}$ $=$ $\sum_{i=0}^{m}{m\choose i}(t;\kappa)_{i}(z;\kappa)_{m-i}\,,$
(3.3) $(t;\kappa)_{i}(t;\kappa)_{j}$ $=$ $\sum_{l=0}^{\min(i,j)}\binom{i}{l}\binom{j}{l}l!\,\kappa^{l}\,(t;\kappa)_{i+j-l}\,,$
(3.4) $(t;\kappa)_{m}$ $=$ $\sum_{l=0}^{m}s_{1}(m,l)\,\kappa^{m-l}t^{l},$
(3.5) $t^{m}$ $=$ $\sum_{l=0}^{m}s_{2}(m,l)\,\kappa^{m-l}(t;\kappa)_{l},$
@@
We call a polynomial $f(t)\in{\mathbb{F}}_{p}[t]$ a quasi-constant if $f(t-\kappa)=f(t)$. The quasi-constants are polynomials in $t^{p}-\kappa^{p-1}t$. A Pochhammer polynomial $(t;\kappa)_{m}$ is a quasi-constant if $m$ is divisible by $p$. Then $(t;\kappa)_{pa}=(t^{p}-\kappa^{p-1}t)^{a}$.
@@
4. $p$-hypergeometric solutions for $\kappa\in{\mathbb{F}}_{p}^{\times}\subset{\mathbb{K}}$
4.1. Solutions in ${\operatorname{Sing}}L^{\otimes n}[n-2]$
In this paper we study solutions of the qKZ equations with values in $V:={\operatorname{Sing}}L^{\otimes n}[n-2]$. The space $L^{\otimes n}[n-2]$ has a basis
$v^{(i)}=v_{1}\otimes\dots\otimes v_{2}\otimes\dots\otimes v_{1}\,,\qquad i=1,\dots,n,$
where the only $v_{2}$ stays at the $i$-th place. In this basis, the subspace $V$ consists of all vectors with the sum of coordinates equal to zero. We identify $L^{\otimes n}[n-2]$ with ${\mathbb{K}}^{n}$ and the subspace $V$ with the vector space
(4.1) $\{x\in{\mathbb{K}}^{n}\mid x_{1}+\dots+x_{n}=0\}.$
4.2. Master polynomial and weight functions
@@
(4.2) $\kappa k\,\equiv\,-1\pmod{p}.$
Define master polynomial
$\Phi(t,z;\kappa)=\prod_{a=1}^{n}(t-z_{a};\kappa)_{k}\,,$
where $k$ is defined in (4.2). For $1\leqslant a\leqslant n$, define the weight functions
(4.3) $\eta_{a}(t,z)$ $=$ $\frac{1}{t-z_{a}}\prod_{j=1}^{a-1}\frac{t-z_{j}+1}{t-z_{j}}\,,$
$Q_{a}(t,z;\kappa)$ $=$ $\Phi(t,z;\kappa)\,\eta_{a}(t,z).$
@@
(4.4) $Q_{a}(t,z;\kappa)=\left(\prod_{j=1}^{a-1}(t-z_{j}-\kappa;\kappa)_{k}\right)(t-z_{a}-\kappa;\kappa)_{k-1}\left(\prod_{j=a+1}^{n}(t-z_{j};\kappa)_{k}\right).$
@@
(4.5) $Q(t,z;\kappa)=(Q_{1}(t,z;\kappa),\dots,Q_{n}(t,z;\kappa))^{\intercal}=\sum_{i}Q^{i}(z;\kappa)(t;\kappa)_{i}\,,$
where $M^{\intercal}$ denotes the transpose matrix of a matrix $M$ and
$Q^{i}(z;\kappa)=(Q^{i}_{1}(z;\kappa),\dots,Q^{i}_{n}(z;\kappa))^{\intercal}$
are vectors of polynomials in $z$.
@@
Theorem 4.3 ([MV, Theorem 5.1]).
For any positive integer $\ell$, the vector $Q^{\ell p-1}(z;\kappa)$ of polynomials in $z$ is a solutions of the qKZequations with step $\kappa$ and values in $V={\operatorname{Sing}}L^{\otimes n}[n-2]$, in particular,
(4.6) $\sum_{a=1}^{n}Q^{\ell p-1}_{a}(z;\kappa)=0,$
see (4.1).
Let $[x]$ denote the integer part of a real number $x$. The vector $Q^{\ell p-1}(z;\kappa)$ is zero if $\ell\not\in\big{\{}1,\dots,\big{[}\frac{n{k}}{p}\big{]}\big{\}}$ for degree reasons. The vectors of polynomials
(4.7) $Q^{\ell p-1}(z;\kappa),\qquad\ell=1,\dots,\left[\frac{n{k}}{p}\right],$
are called the $p$-hypergeometric solutions of the qKZ equations with step $\kappa$ and values in $V$. Denote
(4.8) $d(\kappa)=\left[\frac{n{k}}{p}\right].$
@@
4.3. Step $-\kappa$
The integer $k$ satisfies the inequalities $0<k<p$ and the congruence $\kappa k\equiv-1\pmod{p}$,  see (4.2). Then the integer $p-k$ satisfies the inequalities $0<p-k<p$ and the congruence $-\kappa(p-k)\equiv-1\pmod{p}$. Hence
$\Phi(t,z;-\kappa)$ $=$ $\prod_{a=1}^{n}(t-z_{a};-\kappa)_{p-k}\,.$
Recall that that

### v2

@@ 2.5 KZ equations as a limit of qKZ equations
$f(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})=K_{a}(z;\kappa)f(z),\qquad a=1,\dots,n.$
@@
$g(w_{1},\dots,w_{a}-\alpha\kappa,\dots,w_{n};\alpha)=K_{a}(w/\alpha;\kappa)g(w;\alpha),\qquad a=1,\dots,n,$ (2.7)
@@
$K_{a}(w/\alpha;\kappa)=R^{(a,a-1)}(w_{a}-w_{a-1}-\alpha\kappa;\alpha)\cdots R^{(a,1)}(w_{a}-w_{1}-\alpha\kappa;\alpha)$
[…]
@@
$g-\alpha\kappa\frac{\partial g}{\partial w_{a}}+\mathcal{O}\big(\alpha^{2}\big)=\big(1-\alpha H_{a}(w)+\mathcal{O}\big(\alpha^{2}\big)\big)g.$
In the limit $\alpha\to 0$, we obtain the KZ differential equations
$\kappa\frac{\partial g}{\partial w_{a}}(w)=H_{a}(w)g(w).$
@@
$C_{a}^{\operatorname{KZ}}(z;\kappa)=\lim_{\alpha\to 0}\hat{C}_{a}(z_{1}/\alpha,\dots,z_{n}/\alpha;\kappa)/\alpha^{p}.$
@@
$C_{a}^{\operatorname{KZ}}(w;\kappa)=\lim_{\alpha\to 0}(C_{a,\alpha}(w;\kappa)-1)/\alpha^{p}.$ (2.8)
@@
$(\nabla_{a,\alpha}g)(w)=g-\alpha H_{a}(w)g+\alpha\kappa\frac{\partial g}{\partial w_{a}}+\mathcal{O}\big(\alpha^{2}\big).$
@@
$(\nabla_{a,\alpha}-1)g=\alpha\nabla^{\operatorname{KZ}}_{a}g+\mathcal{O}\big(\alpha^{2}\big).$
@@
$C_{a,\alpha}(w;\kappa)-1=(\nabla_{a,\alpha})^{p}-1=(\nabla_{a,\alpha}-1)^{p}=\alpha^{p}\big(\nabla^{\operatorname{KZ}}_{a}\big)^{p}+\mathcal{O}\big(\alpha^{p+1}\big).$
@@
The degree of the reduced $p$-curvature operator $\tilde{C}_{a}(z;\kappa)$ as a polynomial in $z$ is not greater than (n2)p(n-2)p. Moreover, we have $\tilde{C}_{a}(z;\kappa)=\tilde{C}_{a}^{\operatorname{KZ}}(z;\kappa)+f_{1}(z)$, where $f_{1}(z)$ is a polynomial in $z$ of degree less than (n2)p(n-2)p.
@@
The first statement of the corollary follows from the fact that the denominator of the reduced $p$-curvature operator $\hat{C}_{a}(z;\kappa)$ is of degree (n1)p(n-1)p. Hence the numerator is of degree not greater than (n2)p(n-2)p for the limit in (2.8) to exist. The second statement is clear. ∎
2.6 Limit of solutions
Let $f(z_{1},\dots,z_{n})$ be a polynomial solution of the qKZ equations (2.3). Let $\deg f(z)=d$ and $f(z)=f_{0}(z)+f_{1}(z)$}, where $f_{0}(z)$ is a homogeneous polynomial of degree $d$ and $f_{1}(z)$ is a polynomial of degree less than $d$.
@@ 2.6 Limit of solutions
Let $\alpha$ be a formal parameter. Define $g(w_{1},\dots,w_{n};\alpha)=\alpha^{d}f(w_{1}/\alpha,\dots,w_{n}/\alpha)$. Then $g(w_{1},\dots,w_{n};\alpha)=f_{0}(w)+\alpha g_{1}(w;\alpha)$, where $g_{1}(w;\alpha)=\alpha^{d-1}f_{1}(w/\alpha)$ is a polynomial in $w$ and $\alpha$. We have
$g(w_{1},\dots,w_{a}-\alpha\kappa,\dots,w_{n};\alpha)=K_{a}(w/\alpha;\kappa)g(w;\alpha),\qquad a=1,\dots,n.$
@@
$n_{a}(w;\alpha)=\big(w_{a}-w_{a-1}-\alpha\kappa-\alpha P^{(a,a-1)}\big)\cdots\big(w_{a}-w_{a+1}-\alpha P^{(a,a+1)}\big)$
@@
$d_{a}(w;\alpha)g(w_{1},\dots,w_{a}-\alpha\kappa,\dots,w_{n};\alpha)=n_{a}(w;\alpha)g(w;\alpha)$ (2.9)
in the variables $w$ and $\alpha$. Equation (2.9) gives an equation in $w$ for every fixed power of $\alpha$ in (2.9). The equation corresponding to the first power of $\alpha$ after division by $d_{a}(w;0)$ becomes
$\kappa\frac{\partial f_{0}}{\partial w_{a}}(w)=\sum_{j\neq a}\frac{P^{(a,j)}-1}{w_{a}-w_{j}}f_{0}(w).$
2.7 Dual qKZ equations
Denote $W=L^{\otimes n}$. Let $W^{*}$ be the dual space of $W$, and $\langle,\rangle\colon W^{*}\otimes W\to{\mathbb{K}}$ the canonical pairing. Let $K_{a}^{*}(z;\kappa)\colon W^{*}\to W^{*}$ be the operators dual to the operators $K_{a}(z;\kappa)$. Denote $\tilde{K}_{a}(z;\kappa)=(K_{a}^{*}(z;\kappa))^{-1}$.
@@ 2.7 Dual qKZ equations
$\tilde{K}_{a}(z_{1},\dots,z_{b}-\kappa,\dots,z_{n};\kappa)\tilde{K}_{b}(z;\kappa)=\tilde{K}_{b}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n};\kappa)\tilde{K}_{a}(z;\kappa)$
@@
$\langle x,y\rangle=\big\langle\tilde{K}_{a}(z;\kappa)x,K_{a}(z;\kappa)y\big\rangle$ (2.10)
for all $a$ and $x\in W^{*},y\in W$.
@@
$\tilde{s}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})=\tilde{K}_{a}(z;\kappa)\tilde{s}(z),\qquad a=1,\dots,n,$ (2.11)
@@
$\langle\tilde{s}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n}),s(z_{1},\dots,z_{a}-\kappa,\dots,z_{n},)\rangle=\langle\tilde{s}(z),s(z)\rangle.$
The set of vectors $\{v_{I}\mid I\subset\{1,\dots,n\}\}$ is a basis of $W$. Define the nondegenerate symmetric bilinear form $S$ on $W$ by the formula $S(v_{I},v_{J})=\delta_{I,J}$. The form is called the (tensor) Shapovalov form. We identify $W^{*}$ and $W$ with the help of the Shapovalov form.
@@
$\tilde{K}_{a}(z;\kappa)=R^{(a,a-1)}(-z_{a}+z_{a-1}+\kappa)\cdots R^{(a,1)}(-z_{a}+z_{1}+\kappa)$
[…]
@@
$\tilde{K}_{a}(z_{1},\dots,z_{n};\kappa)=K_{a}(-z_{1},\dots,-z_{n};-\kappa),$
@@
$\tilde{s}(z_{1},\dots,z_{a}-\kappa,\dots,z_{n})=K_{a}(-z_{1},\dots,-z_{n};-\kappa)s(z),\qquad a=1,\dots,n.$
@@
Let $s(z)$ be a solution of the qKZ equations (2.3) with step $\kappa$ and $\tilde{s}(z)$ a solution of the qKZ equations (2.3) with step $\kappa$. Then the function $S(\tilde{s}(-z),s(z))$ is $\kappa$-periodic with respect to every variable $z_{a}$,
$S(\tilde{s}(-z_{1},\dots,-(z_{a}-\kappa),\dots,-z_{n}),s(z_{1},\dots,z_{a}-\kappa,\dots,z_{n}))=S(\tilde{s}(-z),s(z)).$
@@
3 Pochhammer polynomials
For $\kappa\in{\mathbb{K}}$ and $m$ a positive integer, define Pochhammer polynomial $(t;\kappa)_{m}\in{\mathbb{K}}[t]$ by the formula $(t;\kappa)_{m}=\prod_{i=1}^{m}(t-(i-1)\kappa)$}. [^1] See https://en.wikipedia.org/wiki/Falling_and_rising_factorials. We have
@@ 3 Pochhammer polynomials
$(t+z;\kappa)_{m}=\sum_{i=0}^{m}{m\choose i}(t;\kappa)_{i}(z;\kappa)_{m-i},$
$(t;\kappa)_{i}(t;\kappa)_{j}=\sum_{l=0}^{\min(i,j)}\binom{i}{l}\binom{j}{l}l!\kappa^{l}(t;\kappa)_{i+j-l},$ (3.1)
$(t;\kappa)_{m}=\sum_{l=0}^{m}s_{1}(m,l)\kappa^{m-l}t^{l},$
$t^{m}=\sum_{l=0}^{m}s_{2}(m,l)\kappa^{m-l}(t;\kappa)_{l},$ (3.2)
@@
We call a polynomial $f(t)\in{\mathbb{F}}_{p}[t]$ a quasi-constant if $f(t-\kappa)=f(t)$. The quasi-constants are polynomials in $t^{p}-\kappa^{p-1}t$. A Pochhammer polynomial $(t;\kappa)_{m}$ is a quasi-constant if $m$ is divisible by $p$. Then $(t;\kappa)_{pa}=\big(t^{p}-\kappa^{p-1}t\big)^{a}$.
@@
4 $\boldsymbol{p}$-hypergeometric solutions for $\boldsymbol{\kappa\in{\mathbb{F}}_{p}^{\times}\subset{\mathbb{K}}}$
4.1 Solutions in $\boldsymbol{\operatorname{Sing}L^{\otimes n}[n-2]}$
In this paper, we study solutions of the qKZ equations with values in $V:=\operatorname{Sing}L^{\otimes n}[n-2]$. The space $L^{\otimes n}[n-2]$ has a basis $v^{(i)}=v_{1}\otimes\dots\otimes v_{2}\otimes\dots\otimes v_{1}$, $i=1,\dots,n$, where the only $v_{2}$ stays at the $i$-th place. In this basis, the subspace $V$ consists of all vectors with the sum of coordinates equal to zero. We identify $L^{\otimes n}[n-2]$ with ${\mathbb{K}}^{n}$ and the subspace $V$ with the vector space
$\{x\in{\mathbb{K}}^{n}\mid x_{1}+\dots+x_{n}=0\}.$ (4.1)
4.2 Master polynomial and weight functions
@@
$\kappa k\equiv-1\pmod{p}$. (4.2)
Define master polynomial $\Phi(t,z;\kappa)=\prod_{a=1}^{n}(t-z_{a};\kappa)_{k}$, where $k$ is defined in (4.2). For $1\leqslant a\leqslant n$, define the weight functions
$\eta_{a}(t,z)=\frac{1}{t-z_{a}}\prod_{j=1}^{a-1}\frac{t-z_{j}+1}{t-z_{j}},\qquad Q_{a}(t,z;\kappa)=\Phi(t,z;\kappa)\eta_{a}(t,z).$
@@
$Q_{a}(t,z;\kappa)=\left(\prod_{j=1}^{a-1}(t-z_{j}-\kappa;\kappa)_{k}\right)(t-z_{a}-\kappa;\kappa)_{k-1}\left(\prod_{j=a+1}^{n}(t-z_{j};\kappa)_{k}\right).$
@@
$Q(t,z;\kappa)=(Q_{1}(t,z;\kappa),\dots,Q_{n}(t,z;\kappa))^{\intercal}=\sum_{i}Q^{i}(z;\kappa)(t;\kappa)_{i},$ (4.3)
where $M^{\intercal}$ denotes the transpose matrix of a matrix $M$ and $Q^{i}(z;\kappa)=\big(Q^{i}_{1}(z;\kappa),\dots,Q^{i}_{n}(z;\kappa)\big)^{\intercal}$ are vectors of polynomials in $z$.
@@
Theorem 4.3 ([5, Theorem 5.1]).
For any positive integer $\ell$, the vector $Q^{\ell p-1}(z;\kappa)$ of polynomials in $z$ is a solutions of the qKZ equations with step $\kappa$ and values in $V=\operatorname{Sing}L^{\otimes n}[n-2]$, in particular, $\sum_{a=1}^{n}Q^{\ell p-1}_{a}(z;\kappa)=0$, see (4.1).
Let $[x]$ denote the integer part of a real number $x$. The vector $Q^{\ell p-1}(z;\kappa)$ is zero if $\ell\not\in\big\{1,\dots,\big[\frac{n{k}}{p}\big]\big\}$ for degree reasons. The vectors of polynomials $Q^{\ell p-1}(z;\kappa)$, $\ell=1,\dots,\big[\frac{n{k}}{p}\big]$, are called the $p$-hypergeometric solutions of the qKZ equations with step $\kappa$ and values in $V$. Denote $d(\kappa)=\big[\frac{n{k}}{p}\big]$}.
@@
4.3 Step $\boldsymbol{-\kappa}$
The integer $k$ satisfies the inequalities $0<k<p$ and the congruence $\kappa k\equiv-1\pmod{p}$, see (4.2). Then the integer pkp-k satisfies the inequalities 0<pk<p0<p-k<p and the congruence $\kappa(p-k)\equiv-1\pmod{p}$. Hence $\Phi(t,z;-\kappa)=\prod_{a=1}^{n}(t-z_{a};-\kappa)_{p-k}$. Recall that